OpenStudy (anonymous):

Which is the simplified form of 12 times f to the sixth power times g all over 16 times g squared.? A:f to the sixth power over 4 times g squared. B:3 times f to the sixth power all over 4 times g squared. C:3 times f to the sixth power all over 4 times g. D:f to the sixth power over 4 times g.

OpenStudy (anonymous):

\[\frac{12f^6g}{16g^2}\]

OpenStudy (anonymous):

\[\frac{12}{16}=\frac{3}{4}\] and \[\frac{f^6g}{g^2}=\frac{f^6}{g}\] so your answer should be \[\frac{3f^6}{4g}\]

OpenStudy (anonymous):

as usual, it is C it is almost always C